11. A Fahrenheit and a Celsius thermometer are both immersed in a fluid. a) If the two numerical readings

are identical, what is the temperature of the fluid expressed in o K and o R. b) If the Fahrenheit

temperature is numerically twice that of the Celsius reading, what is the temperature of the fluid

expressed in o K and o R?


1
Expert's answer
2021-09-12T00:30:29-0400

Fahrenheit and Celsius degrees can be interconverted through the formulae

"\degree F=\frac{9}{5}(\degree C)+32"

a)

To have the same value;

Let °C=x, then °F=x

"x=\frac{9}{5}x+32"

"x-\frac{9}{5}x=32"

"-\frac{4}{5}x=32"

-4x=32*5

x"=\frac{160}{-4}"

x=-40

Thus, -40°C=-40°F

°K=°C+273.15

=-40+273.15

=233.15°K

°R=°F+459.67

=-40+459.67

=419.67°R

b)

°F is twice °C

Let °C=x, "\therefore" °F=2x

2x="\frac{9}{5}x+32"

"2x-\frac{9}{5}x=32"

"\frac{1}{5}x=32"

x=32*5=160

Thus 160°C=320°F

°K=°C+273.15

=160+273.15

=433.15°K

°R=°F+459.67

=320+459.67

=779.67°R